OpenStudy (anonymous):

How do you graph this x2- 3x + 2<= 0

OpenStudy (anonymous):

\[x^2+3x+2\le0\]

OpenStudy (anonymous):

wow!!!! i wouldnt know sorry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

k. but anyways i think your cute tho.

OpenStudy (helder_edwin):

first solve it as if it were an equation \[ \large x^2-3x+2=0 \]

OpenStudy (anonymous):

it's a parabola. To find where it crosses x-axis, write it like this: x^2-3x+2=(x-1)(x-2)=y so it will cross x-axis at x=1 and x=2 since you lookiing for points that are less than 0 it will be the part that is between this two points: